OpenStudy (anonymous):

In five years a boy will be two-third as old as his uncle. Three years ago he was half as old as the uncle is now. how old are the boy and his uncle?

OpenStudy (triciaal):

let the boys age = b and the uncle's age = u in 5 years b + 5 = 2/3 * u 3 years ago b - 3 = 1/2 * u do you see that 1/2* u + 3 = 2/3 * u -5? solve for u then substitute in one of the original equations to find b